OpenStudy (s3a):

Integrating (-2x)/(x^4 + 1) dx. Could someone please help me do this one?

OpenStudy (anonymous):

try \(u=x^2\) and get a familiar anti derivative, namely \(\tan^{-1}(u)\)

OpenStudy (anonymous):

\[u=x^2, du=2xdx,=du=2xdx\] gives you \[-\int\frac{du}{1+u^2}\]

OpenStudy (anonymous):

typo there, i meant \(-du=2xdx\) but the idea is right

OpenStudy (s3a):

Oh, man! That was easy!

OpenStudy (anonymous):

yup

OpenStudy (s3a):

Thanks! :D

OpenStudy (anonymous):

yw